Required information Exercise 3-31 and 3-32 (Algo) (LO 3-1) (The following information applies to the questions displayed belRequired A Required B Require What will be the operating profit? Operating profitRequired A Required B Required C Required D What is the impact on operating profit if the sales price decreases by 15 percentRequired A Required B Required C Required D What is the impact on operating profit if variable costs per unit decrease by 15Required A Required B Required C Required D Suppose that fixed costs for the year are 15 percent lower than projected, and va
Answer #1

1) compute operating profit

Particular per unit amount ($)
Sale(6000×12)72000 $16 1152000
(-) variable expenses $4 (288000)
Contribution 864000
(-) fixed expenses (612000)
operating profit 252000

2)if sale price decreases by 15%

   $16-15%=13.6, new selling price= 13.6

If sales price decreases but variable cost per Will not change and remains constant therefore

Particular per unit amount ($)
Sale (6000×12)72000 13.6 979200
(-) variable expenses 4 (288000)
Contribution 691200
(-) fixed expenses (612000)
Operating income 79200

if sales price increase by 30%

   $16+30% =$20.8, new selling price= $20.8

Particular per unit amount ($)
Sales (6000×12)72000 $20.8 1497600
(-) variable expenses $4 (288000)
Contribution 1209600
(-) fixed expenses (612000)
operating income 597600

3)if variable cost per unit decrease by 15%

   $4-15%=$3.4, New variable price per unit=$3.4

Particular per unit amount ($)
Sales (6000×12)72000 $16 1152000
(-) variable cost $3.4 (244800)
Contribution 907200
(-) fixed expenses (612000)
Operating income 295200

if variable cost per unit increase by 30%

  $4+30%=$5.2, then New variable cost per unit=$5.2

particular per unit amount ($)
Sales (6000×12)72000 $16 1152000
(-) variable expenses $5.2 (374400)
Contribution 777600
(-) fixed cost (612000)
Operating income 165600

3) if Fixed costs decrease by 15% , variable cost per unit increase by 15%

New fixed cost= $51000×12-15%=$520200

New variable cost per unit=$4+15%=$4.6

Particular per unit amount ($)
Sales(6000×12)72000 $16 1152000
(-) variable expenses $4.6 (331200)
Contribution 820800
(-) fixed expenses (520200)
Operating income 300600

Therefore profit had increase

$300600 - $252000= $48600

   ALL THE BEST

   PLEASE DO SUPPORT US

   ANY DOUBT PLEASE COMMENT BELOW

THANK YOU